Jason emptied his piggy bank and counted exactly $10 in nickels, dimes and quarters There were twice as many dimes as nickels. There were twice as many quarters as dimes. How many nickels did he have?
Answer by stanbon(75887)

Jason emptied his piggy bank and counted exactly $10 in nickels, dimes and quarters There were twice as many dimes as nickels. There were twice as many quarters as dimes. How many nickels did he have?--------------------
-------
5n + 10d + 25q = 1000 cents
d = 2n
q = 2d
--------
Substitute for "n" and "q" and solve for "d"::
5(d/2) + 10d + 25(2d) = 1000
5d + 20d + 100d = 2000
125d = 2000
d = 16 (# of dimes)
--------------------------
n = d/2 = 8 (# of nickels)
------
q = 2d = 32 (# of quarters)
--------------
Cheers,
Stan H.
--------------